Session: Feb 2000 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

HartRAO (Jonathan Quick at Sun Feb 13 15:00:20 MET 2000) Minor failures
Lost the few minutes due to local software bug. Lost final scan to RTC testing induced PCFS failure in preparation for RTC001.

Medicina (Zacchiroli at Sun Feb 13 15:59:41 MET 2000) Success
no known problems

Torun (K.Borkowski at Sun Feb 13 17:07:52 MET 2000) Success
No known problems.

Onsala (A. Polatidis at Mon Feb 14 10:16:07 MET 2000) Success

Westerbork (Tony Foley at Mon Feb 14 10:50:12 MET 2000) Success
Used 13 telescopes in tied array (RT8 not working). No other known problems

Effelsberg (A. Kraus at Tue Feb 15 08:44:15 MET 2000) Success

Jodrell (Alastair Gunn at Wed Feb 16 05:47:06 MET 2000) Success
Tsys data not reliable due to GPIB errors. No other known problems.

Noto (C.Trigilio at Thu Feb 17 12:35:36 MET 2000) Success

Onsala (A. Polatidis at Mon Feb 21 13:58:34 MET 2000) Minor failures
NOTE (21 Feb 2000): There was a ~200 nsec Formatter clock jump which occured between Day 42 and Day 44, before the start of this experiment. The Formatter was re-synched on Day 52 . See the gps-fmout numbers on more information.

Calibration (Alastair Gunn at Wed Apr 19 12:59:15 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/gp025a.antab for details